# InceptionMamba: An Efficient Hybrid Network with Large Band Convolution and Bottleneck Mamba

### v2

@@ Abstract
Within the family of convolutional neural networks, InceptionNeXt has shown excellent competitiveness in image classification and a number of downstream tasks. Built on parallel one-dimensional strip convolutions, however, it suffers from limited ability of capturing spatial dependencies along different dimensions and fails to fully explore spatial modeling in local neighborhood. Besides, inherent locality constraints of convolution operations are detrimental to effective global context modeling. To overcome these limitations, we propose a novel backbone architecture termed InceptionMamba in this study. More specifically, the traditional one-dimensional strip convolutions are replaced by orthogonal band convolutions in our InceptionMamba to achieve cohesive spatial modeling. Furthermore, global contextual modeling can be achieved via a bottleneck Mamba module, facilitating enhanced cross-channel information fusion and enlarged receptive field. Extensive evaluations on classification and various downstream tasks demonstrate that the proposed InceptionMamba achieves state-of-the-art performance with superior parameter and computational efficiency. The source code will be available at https://github.com/Wake1021/InceptionMamba.
@@ 1 Introduction
In Convolutional Neural Networks (CNNs) [1, 2, 3], the success of Inception architecture is largely attributed to efficient strip convolution, evolving from local feature extraction to dynamic parameter optimization [3]. Originated from InceptionNeXt [4], it is beneficial for reducing the number of parameters through asymmetric convolution decomposition (such as splitting a 7×7 convolution into 1×7 and 7×1 strip convolutions). As a representative member within Inception family, InceptionNeXt has showcased the efficacy of strip convolutions in balancing computational efficiency and representational abilities. In InceptionNeXt, large-kernel depthwise convolutions are decomposed into four efficient branches including parallel one-dimensional strip convolutions, such that large-kernel CNNs are accelerated with maintained performance.
Despite achieving huge success, conventional one-dimensional strip convolutions in InceptionNeXt suffer twofold inherent limitations, which hinder their potential in complex visual understanding scenarios. On the one hand, the one-dimensional strip convolutions are independently computed in separate channel groups, lacking sufficient spatial modeling capacity in capturing complex visual contents. For instance, while horizontal kernels effectively characterize axis-aligned patterns, their weakened responses to orthogonal structures (e.g., vertical edges) results in less discriminative encoding. Additionally, the local inductive bias of CNNs, arising from utilizing limited-size convolutional kernels (e.g., 3$\times$3 convolutions) for hierarchical local feature extraction, fundamentally constrains their capacity in global contextual relationship modeling, thereby suffering from restricted receptive field. While recent efforts to addressing this limitation have explored oversized kernels [5] (e.g., 31$\times$31 convolutions) and dilated convolutions [6] to expand receptive fields, these approaches are challenged by accuracy-efficiency trade-off: enlarged kernels introduce a quadratic increase in computational complexity $(O(k^{2}))$, while dilated convolutions are prone to grid artifacts that degrade spatial coherence in feature representations.
@@
To overcome the aforementioned drawbacks, we propose an efficient hybrid backbone architecture with large band convolutions and bottleneck Mamba termed InceptionMamba. Inherited from Inception architecture, our InceptionMamba maintains Inception-style depthwise convolutions with efficient parallel structures, yet features carefully designed orthogonal band convolutions instead of the traditional one-dimensional stripe convolutions. Thus, multi-dimensional visual cues can be discovered for cohesive spatial modeling, leading to comprehensive embedding with improved representational capacity. In addition, to improve inter-channel interaction ability, we incorporate a bottleneck Mamba module into the token mixture for facilitating cross-channel information integration and enlarged receptive field. The advantages of our framework are illustrated in Fig. 2, demonstrating that InceptionMamba can capture semantically correlated local neighborhood with sufficient discriminative power. For instance, our method can accurately focus on object-aware regions surrounding panda eye and body, whereas InceptionNeXt only identify isolated eye-centric areas with degraded spatial modeling capacity.
@@
    In contrast to one-dimensional strip convolution involved in InceptionNeXt [4], our carefully designed orthogonal band convolutions significantly contribute to characterizing semantically cohesive visual cues in local adjacent neighborhood, thereby demonstrating improved spatial modeling capacity.
@@ 2.1 Inception Architecture
The evolution of the Inception architecture began with the core idea of multi-scale feature fusion. Earliest Inception architecture [10] introduces parallel multi-branch convolutions (1$\times$1, 3$\times$3, 5$\times$5, and pooling) along with 1$\times$1 dimensionality reduction to overcome the single-scale limitation. Subsequent improvements of Inception-v2 [11] and Inception-v3 [11] are manifested in enhancing computational efficiency via convolution decomposition (e.g., splitting one 5$\times$5 convolution into two 3×3 convolutions) and batch normalization (BN). Similar pure Inception variant is Inception-v4 [12] which contains multiple consecutive Inception blocks with parallel structures. More advanced Inception-ResNet is a hybrid Inception network which integrates the residual learning into Inception architecture, achieving network training speedup through skip connections. A milestone work within Inception family emerges as InceptionNeXt [4] which decomposes large-kernel depthwise convolutions into multiple parallel branches of smaller kernels (e.g., square 3$\times$3 convolutions and one-dimensional strip convolutions), achieving desirable accuracy-throughputs trade-off in a series of downstream tasks. It is significantly beneficial for accelerating large-kernel networks with maintained performance.
@@ 2.2 Mamba
To achieve efficient discrete computation, Mamba [13, 14] introduces zero-order preserving discretization method, which converts the continuous system into a recursive form through the time scale parameter $\Delta$ :
@@
Recent advancements in visual Mamba models demonstrate the potential of SSM in a variety of vision tasks. In Vision Mamba [15], a bidirectional Mamba block was proposed for efficient high-resolution image processing. VMamba [9] introduces Cross-Scan strategy to strengthen global context modeling. MambaVision [8] combines linear complexity of Mamba with Transformer self-attention in a hybrid architecture to enhance efficiency and global spatial reasoning. Recent research explores the essence of Mamba, and reveals that Mamba is not ideally suited for all vision tasks. Meanwhile, a range of MambaOut [16] models are obtained by stacking Mamba blocks while removing their core token mixer SSM.
@@ 3 Method
Encouraged by recent success of vision Mamba models, we propose to deeply integrate Mamba SSM [13, 14] into efficient InceptionNeXt [4] architecture with parallel structures. The former, which enjoys linear-complexity sequence modeling capability, can capture long-range spatial dependencies, whilst the latter excels at extracting key features while maintaining network efficiency. In this sense, our synergistic design helps to preserve strengths of CNNs in efficient local reasoning while enhancing model adaptability to complex scenes through global dynamic feature selection mechanism of Mamba [13, 14], establishing a new foundation paradigm for a series of downstream vision tasks.
@@ 3.1 Overview of InceptionMamba
Inspired by InceptionNeXt [4], we have designed a four-stage InceptionMamba architecture as illustrated in Fig. 3. For a given input tensor $X\in\mathbb{R}^{B\times C\times H\times W}$ where $B$, $C$, $H$, $W$ represent the batch size, number of channels, height, and width, respectively, the feature processing begins with a Local Mixer module that emphasizes local spatial information encoding similar to token mixer in InceptionNeXt [4]:
$X^{{}^{\prime}}=\text{LocalMixer}(X)$ (5)
@@
As the cornerstone of our method, the core InceptionMamba block constitute four-stage framework resembling ConvNeXt [3] and InceptionNeXt [4], leading to a family of InceptionMamba model with varying sizes. More specifically, the block numbers in the three model variants are [3, 3, 12, 3], [4, 4, 32, 4] and [4, 4, 34, 4] for the Tiny (T), Small (S) and Base (B), respectively. Starting from an input image of $3\times H\times W$, our model first applies patch embedding to process the visual data. In each stage, subsequently, downsampling operation is combined with multiple InceptionMamba blocks to progressively reduce spatial resolution, i.e., spatial size $H\times W$ is reduced by 4 times, 16 times, 32 times, and 64 times respectively, while expanding the channel dimension (C). This pyramid-like hierarchical architecture ends with global average pooling and classifiers for the final prediction. The architecture configurations of the three InceptionMamba variants can be referred to Table 1.
@@ 4 $\frac{H}{32}\times\frac{W}{32}$ DownSampling Kernel Size $3\times 3$, stride 2
Figure 4: Illustrative comparison of ConvNeXt [3], InceptionNeXt [4], and our InceptionMamba block. InceptionMamba consists of two key components, namely LocalMixer and GlobalMixer. The former inherits efficient parallel structure from InceptionNeXt [4] to accelerate large-kernel depthwise convolutions. Different from InceptionNeXt [4] using one-dimensional strip convolutions, our InceptionMamba takes advantage of orthogonal band convolutions to capture multi-dimensional visual cue for cohesive feature embedding. Furthermore, the GlobalMixer adopts a carefully designed bottleneck Mamba [13, 14] to facilitate inter-channel information with enlarged receptive field. Compared to ConvNeXt [3] and InceptionNeXt [4], our InceptionMamba enjoys efficient local spatial modeling capacity while excels in capturing long-range dependencies.
3.2 LocalMixer with large-kernel band convolutions
Inspired by SLaK [17] and InceptionNeXt [4], we devise a multi-branch structure in our LocalMixer to address the limitation of one-way large kernels. Inherited from InceptionNeXt[4] block, the parallel structure allows speedup of large-kernel depthwise convolutions for efficient local spatial modeling. Different from its predecessor, however, our LocalMixer employs orthogonal large-kernel band convolutions instead of one-dimensional strip convolutions in InceptionNeXt [4]. Band convolution can focus on larger areas with improved local modeling capability. This advantage can be intuitively shown in Fig. 2, clearly demonstrating that our InceptionMamba generates more spatially cohesive activations on the object-aware regions compared to InceptionNeXt[4].
Mathematically, the input features $X$ are divided into three groups along the channel dimension via our LocalMixer as follows:
@@ 3.2 LocalMixer with large-kernel band convolutions
Notably, identity mapping operations are applied to the majority of channel groups in the LocalMixer while depthwise convolutions are only imposed on a small subset of groups. This is consist with InceptionNeXt [4] which reveals that depthwise convolutions are not required for most channels.
@@ 3.3 GlobalMixer with bottleneck Mamba
After the LocalMixer, we design a GlobalMixer module for inter-channel interaction at the Token layer. Traditional methods including ConvNeXt [3] and InceptionNeXt [4] only pass the concatenated features through a straightforward MLP, leading to inadequate cross-channel interactions due to inherent single-channel processing characteristic of depthwise convolution. To mitigates this drawback, a bottleneck structure is introduced into the GlobalMixer, making use of $1\times 1$ convolution to compress and expand feature channels. This is beneficial for efficient cross-channel fusion in a low-dimensional space while retaining the key information of the original features.
To strengthen global modeling capacities while optimizing computational efficiency, a state-space module (SS2D) [9] is integrated within the bottleneck architecture. This co-design leverages channel compression mechanism (ratio $r=2$) of the bottleneck for significantly reducing the SS2D [9] dimensionality. More importantly, the SS2D [9] module can capture long-range dependencies with the help cross-scan mechanism, while the compressed feature space enjoys desirable parameter efficiency. Mathematically, given the Input $X^{{}^{\prime}}$, our GlobalMixer with a bottleneck Mamba [13, 14] is formulated as follows:
@@ 3.4 Block Comparisons
Since our InceptionMamba is closely related to ConvNeXt [3] and InceptionNeXt [4] frameworks, we compare different building blocks comprising the three frameworks as illustrated in Fig. 4. With the help of the LocalMixer, our InceptionMamba block allows efficient computation of large-kernel depthwise convolutions via a parallel multi-branch structure inherited from the InceptionNeXt [4]. Different from the predecessor which makes use of one-dimensional strip convolutions, our method features orthogonal band convolutions with improved local modeling capacity in characterizing multi-dimensional visual cues. Additionally, the GlobalMixer module with bottleneck Mamba [13, 14] structure can capture long-range dependencies through cross-channel interaction, which is conducive to boosting global modeling capability with expanded receptive field. In contrast, both ConvNeXt [3] and InceptionNeXt [4] demonstrate limited global modeling capacity, lacking GlobalMixer-like component for sufficient information fusion.

### v3

@@ Abstract
Within the family of convolutional neural networks, InceptionNeXt has shown excellent competitiveness in image classification and a number of downstream tasks. Built on parallel one-dimensional strip convolutions, however, it suffers from limited ability of capturing spatial dependencies along different dimensions and fails to fully explore spatial modeling in local neighborhood. Additionally, although InceptionNeXt excels at multi-scale spatial modeling capacity, convolution operations are prone to locality constraints and inherently lack the global receptive field, which is detrimental to effective global context modeling. Recent research demonstrates that Mamba based on state-space models can capture long-range dependencies while enjoying linear computational complexity. Thus, it is necessary to take advantage of the Mamba architecture to improve the long-range modeling capabilities of the InceptionNeXt while maintaining desirable efficiency. Towards this end, we propose a novel backbone architecture termed InceptionMamba in this study. More specifically, the traditional one-dimensional strip convolutions are replaced by orthogonal band convolutions in our InceptionMamba to achieve cohesive spatial modeling. Furthermore, global contextual modeling can be achieved via a bottleneck Mamba module, facilitating enhanced cross-channel information fusion and enlarged receptive field. Extensive evaluations on classification and various downstream tasks demonstrate that the proposed InceptionMamba achieves state-of-the-art performance with superior parameter and computational efficiency. The source code will be available at https://github.com/Wake1021/InceptionMamba.
keywords:
InceptionMamba , Orthogonal Band Convolutions , Global Contextual Modeling , Bottleneck Mamba , Cross-channel Information Fusion
@@ 1 Introduction
In Convolutional Neural Networks (CNNs) [1, 2, 3, 4, 5], the success of Inception architecture is largely attributed to efficient strip convolution, evolving from local feature extraction to dynamic parameter optimization [3]. Originated from InceptionNeXt [6], it is beneficial for reducing the number of parameters through asymmetric convolution decomposition (such as splitting a 7×7 convolution into 1×7 and 7×1 strip convolutions). As a representative member within Inception family, InceptionNeXt has showcased the efficacy of strip convolutions in balancing computational efficiency and representational abilities. In InceptionNeXt, large-kernel depthwise convolutions are decomposed into four efficient branches including parallel one-dimensional strip convolutions, such that large-kernel CNNs are accelerated with maintained performance.
Despite achieving huge success, conventional one-dimensional strip convolutions in InceptionNeXt suffer twofold inherent limitations, which hinder their potential in complex visual understanding scenarios. On the one hand, the one-dimensional strip convolutions are independently computed in separate channel groups, lacking sufficient spatial modeling capacity in capturing complex visual contents. For instance, while horizontal kernels effectively characterize axis-aligned patterns, their weakened responses to orthogonal structures (e.g., vertical edges) results in less discriminative encoding. Additionally, the local inductive bias of CNNs, arising from utilizing limited-size convolutional kernels (e.g., 3$\times$3 convolutions) for hierarchical local feature extraction, fundamentally constrains their capacity in global contextual relationship modeling, thereby suffering from restricted receptive field. While recent efforts to addressing this limitation have explored oversized kernels [7] and dilated convolutions [8] to expand receptive fields, these approaches are challenged by accuracy-efficiency trade-off: enlarged kernels introduce a quadratic increase in computational complexity $(O(k^{2}))$, while dilated convolutions are prone to grid artifacts that degrade spatial coherence in feature representations. In a nutshell, CNNs like InceptionNeXt with local receptive fields and weight-sharing mechanisms are highly efficient at capturing local cues but still struggle to model long-range dependencies. Consequently, it is necessary to improve the long-range modeling capacity of the InceptionNeXt while maintaining efficient parallel structure for enlarged receptive field.
Recently, a new architecture, i.e., Vision Transformer (ViT) [9, 10, 11, 12, 13, 14], has received extensive attention and emerged as a promising alternative to CNNs. ViT leverages self-attention mechanisms to enable global interactions, yet it incurs quadratic computational complexity, which poses enormous challenges to computationally intensive scenarios, e.g., dense prediction tasks. Similar to ViT, the Mamba architecture based on State Space Models has appealing long-range modeling capacity superior to CNNs. It employs a selective state transition mechanism that combines the linear complexity of convolutions with data-dependent global context modeling. Additionally, the gating mechanism allows the Mamba to dynamically focus on key spatial positions while preserving sub-quadratic computational complexity. This attribute well addresses the spatial modeling limitations of the InceptionNeXt, while hardly incurring expensive computational costs like ViT.
@@
Taking advantage of global contextual modeling capability of Mamba, we propose an efficient hybrid backbone architecture with large band convolutions and bottleneck Mamba termed InceptionMamba in this study. Inherited from Inception architecture, our InceptionMamba maintains Inception-style depthwise convolutions with efficient parallel structures, yet features carefully designed orthogonal band convolutions instead of the traditional one-dimensional stripe convolutions. Thus, multi-dimensional visual cues can be discovered for cohesive spatial modeling, leading to comprehensive embedding with improved representational capacity. In addition, to improve inter-channel interaction ability, we incorporate a bottleneck Mamba module into the token mixture for facilitating cross-channel information integration and enlarged receptive field. The advantages of our framework are illustrated in Fig. 2, demonstrating that InceptionMamba can capture semantically correlated local neighborhood with sufficient discriminative power. For instance, our method can accurately focus on object-aware regions surrounding panda eye and body, whereas InceptionNeXt only identify isolated eye-centric areas with degraded spatial modeling capacity.
@@
    In contrast to one-dimensional strip convolution involved in InceptionNeXt, our carefully designed orthogonal band convolutions significantly contribute to characterizing semantically cohesive visual cues in local adjacent neighborhood, thereby demonstrating improved spatial modeling capacity.
@@ 2.1 Inception Architecture
The evolution of the Inception architecture began with the core idea of multi-scale feature fusion. Earliest Inception architecture [18, 19] introduces parallel multi-branch convolutions (1$\times$1, 3$\times$3, 5$\times$5, and pooling) along with 1$\times$1 dimensionality reduction to overcome the single-scale limitation. Subsequent improvements of Inception-v2 [20] and Inception-v3 [20] are manifested in enhancing computational efficiency via convolution decomposition (e.g., splitting one 5$\times$5 convolution into two 3×3 convolutions) and batch normalization (BN). Similar pure Inception variant is Inception-v4 [21] which contains multiple consecutive Inception blocks with parallel structures. More advanced Inception-ResNet [22, 6] is a hybrid Inception network which integrates the residual learning into Inception architecture, achieving network training speedup through skip connections. A milestone work within Inception family emerges as InceptionNeXt [6] which decomposes large-kernel depthwise convolutions into multiple parallel branches of smaller kernels (e.g., square 3$\times$3 convolutions and one-dimensional strip convolutions), achieving desirable accuracy-throughputs trade-off in a series of downstream tasks. It is significantly beneficial for accelerating large-kernel networks with maintained performance.
@@ 2.2 Mamba
To achieve efficient discrete computation, Mamba introduces zero-order preserving discretization method, which converts the continuous system into a recursive form through the time scale parameter $\Delta$ :
@@
Recent advancements in visual Mamba models demonstrate the potential of SSM in a variety of vision tasks. In Vision Mamba [25], a bidirectional Mamba block was proposed for efficient high-resolution image processing. VMamba introduces Cross-Scan strategy to strengthen global context modeling. MambaVision combines linear complexity of Mamba with Transformer self-attention in a hybrid architecture to enhance efficiency and global spatial reasoning. Recent research explores the essence of Mamba, and reveals that Mamba is not ideally suited for all vision tasks. Meanwhile, a range of MambaOut [26] models are obtained by stacking Mamba blocks while removing their core token mixer SSM.
@@ 3 Method
Encouraged by recent success of vision Mamba models, we propose to deeply integrate Mamba SSM into efficient InceptionNeXt architecture with parallel structures. The former, which enjoys linear-complexity sequence modeling capability, can capture long-range spatial dependencies, whilst the latter excels at extracting key features while maintaining network efficiency. In this sense, our synergistic design helps to preserve strengths of CNNs in efficient local reasoning while enhancing model adaptability to complex scenes through global dynamic feature selection mechanism of Mamba, establishing a new foundation paradigm for a series of downstream vision tasks.
@@ 3.1 Overview of InceptionMamba
Inspired by InceptionNeXt, we have designed a four-stage InceptionMamba architecture as illustrated in Fig. 3. For a given input tensor $X\in\mathbb{R}^{B\times C\times H\times W}$ where $B$, $C$, $H$, $W$ represent the batch size, number of channels, height, and width, respectively, the feature processing begins with a ConvMixer module that emphasizes local spatial information encoding similar to token mixer in InceptionNeXt:
$X^{{}^{\prime}}=\text{ConvMixer}(X)$ (5)
@@
As the cornerstone of our method, the core InceptionMamba block constitute four-stage framework resembling ConvNeXt and InceptionNeXt, leading to a family of InceptionMamba model with varying sizes. More specifically, the block numbers in the three model variants are [3, 3, 12, 3], [4, 4, 32, 4] and [4, 4, 34, 4] for the Tiny (T), Small (S) and Base (B), respectively. Starting from an input image of $3\times H\times W$, our model first applies patch embedding to process the visual data. In each stage, subsequently, downsampling operation is combined with multiple InceptionMamba blocks to progressively reduce spatial resolution, i.e., spatial size $H\times W$ is reduced by 4 times, 16 times, 32 times, and 64 times respectively, while expanding the channel dimension (C). This pyramid-like hierarchical architecture ends with global average pooling and classifiers for the final prediction. The architecture configurations of the three InceptionMamba variants can be referred to Table 1.
@@ 4 $\frac{H}{32}\times\frac{W}{32}$ DownSampling Kernel Size $3\times 3$, stride 2
Figure 4: Illustrative comparison of ConvNeXt, InceptionNeXt, and our InceptionMamba block. InceptionMamba consists of two key components, namely ConvMixer and GlobalMixer. The former inherits efficient parallel structure from InceptionNeXt to accelerate large-kernel depthwise convolutions. Different from InceptionNeXt using one-dimensional strip convolutions, our InceptionMamba takes advantage of orthogonal band convolutions to capture multi-dimensional visual cue for cohesive feature embedding. Furthermore, the GlobalMixer adopts a carefully designed bottleneck Mamba to facilitate inter-channel information with enlarged receptive field. Compared to ConvNeXt and InceptionNeXt, our InceptionMamba enjoys efficient local spatial modeling capacity while excels in capturing long-range dependencies.
3.2 ConvMixer with large-kernel band convolutions
Inspired by SLaK [27] and InceptionNeXt, we devise a multi-branch structure in our ConvMixer to address the limitation of one-way large kernels. Inherited from InceptionNeXt block, the parallel structure allows speedup of large-kernel depthwise convolutions for efficient local spatial modeling. Different from its predecessor, however, our ConvMixer employs orthogonal large-kernel band convolutions instead of one-dimensional strip convolutions in InceptionNeXt. Band convolution can focus on larger areas with improved local modeling capability. This advantage can be intuitively shown in Fig. 2, clearly demonstrating that our InceptionMamba generates more spatially cohesive activations on the object-aware regions compared to InceptionNeXt.
Mathematically, the input features $X$ are divided into three groups along the channel dimension via our ConvMixer as follows:
@@ 3.2 ConvMixer with large-kernel band convolutions
Notably, identity mapping operations are applied to the majority of channel groups in the ConvMixer while depthwise convolutions are only imposed on a small subset of groups. This is consist with InceptionNeXt which reveals that depthwise convolutions are not required for most channels.
@@ 3.3 GlobalMixer with bottleneck Mamba
In the traditional InceptionNeXt architecture, the application of depthwise separable convolutions and identity mappings leads to limited channel interaction without sufficiently exploring long-range dependency. Moreover, existing deep architectures commonly suffer from redundant channel information as illustrated in Fig. 5, which inevitably increases the model complexity. To address this limitation, we design a Bottleneck Mamba structure to enhance effective cross-channel interaction and reduce channel redundancy, thereby enhancing the model capacity while hardly affecting the model efficiency. Towards this end, we design a GlobalMixer module for inter-channel interaction at the Token layer after the ConvMixer module. Traditional methods including ConvNeXt and InceptionNeXt only pass the concatenated features through a straightforward MLP, leading to inadequate cross-channel interactions due to inherent single-channel processing characteristic of depthwise convolution. To mitigates this drawback, a bottleneck structure is introduced into the GlobalMixer, making use of $1\times 1$ convolution to compress and expand feature channels. This is beneficial for efficient cross-channel fusion in a low-dimensional space while retaining the key information of the original features.
Figure 5: Visualized feature maps along consecutive channels in an intermediate layer of a pretrained VMamba model. The top-left image shows the input. Substantial redundancies in channel information can be clearly observed.
To strengthen global modeling capacities while optimizing computational efficiency, a state-space module (SS2D)is integrated within the bottleneck architecture. This co-design leverages channel compression mechanism (ratio $r=2$) of the bottleneck for significantly reducing the SS2D dimensionality. More importantly, the SS2D module can capture long-range dependencies with the help cross-scan mechanism, while the compressed feature space enjoys desirable parameter efficiency. Mathematically, given the Input $X^{{}^{\prime}}$, our GlobalMixer with a bottleneck Mamba is formulated as follows:
@@ 3.4 Block Comparisons
Since our InceptionMamba is closely related to ConvNeXt and InceptionNeXt frameworks, we compare different building blocks comprising the three frameworks as illustrated in Fig. 4. With the help of the ConvMixer, our InceptionMamba block allows efficient computation of large-kernel depthwise convolutions via a parallel multi-branch structure inherited from the InceptionNeXt. Different from the predecessor which makes use of one-dimensional strip convolutions, our method features orthogonal band convolutions with improved local modeling capacity in characterizing multi-dimensional visual cues. Additionally, the GlobalMixer module with bottleneck Mamba structure can capture long-range dependencies through cross-channel interaction, which is conducive to boosting global modeling capability with expanded receptive field. In contrast, both ConvNeXt and InceptionNeXt demonstrate limited global modeling capacity, lacking GlobalMixer-like component for sufficient information fusion.